OpenStudy (anonymous):

Find the ratio of the figures below! Please help, I will fan and give a medal to the best/right answer

OpenStudy (anonymous):



OpenStudy (anonymous):

@amistre64 could you help with this?!

OpenStudy (anonymous):

So ratio of what? areas? sides? volumes?

OpenStudy (anonymous):

@luckythebest the areas

OpenStudy (anonymous):

ok so is it given that the triangles are similar?

OpenStudy (anonymous):

sorry I thought I said that, I guess I didn't lol I've been working on school work all day :/

OpenStudy (anonymous):

yes @luckythebest

OpenStudy (anonymous):

Ok. so we can see that 10 units and 8 units sides are corresponding sides. There is one concept. "Ratio of areas of 2 similar triangles is the ratio of SQUARES of corresponding sides." So Ar(big)/Ar(small) = 10^2/8^2 if you can find the value. cheers! :)

OpenStudy (anonymous):

So would I write it as 10^2/8^2 or would I actually do the math so it would be 100/64?! Lol, I'm kind of confused...

OpenStudy (anonymous):

@luckythebest

OpenStudy (anonymous):

no lol its directly 100/64 = 25/16. You want i can prove the above concept here itself.

OpenStudy (anonymous):

but yeah u need to do that math of writing it as 100/64 = 25/16. Ratio hence = 25:16.

OpenStudy (anonymous):

I think I understand it now, thank you @luckythebest

OpenStudy (anonymous):

@HeyImKae but be very careful, that concept applies only to similar triangles, and only the corresponding sides. dont do it for any random non-similar triangles.